# Defeaturing.bib

### v1

@@ 4 A posteriori error analysis
$\sum_{K\in\mathcal{T}_{h}^{0}}||\nabla(v_{0}-I_{h}v_{0})||_{K}^{2}\lesssim||\nabla v_{0}||_{\Omega_{0}}^{2},$ (49)
Since ${I_{h}v_{0}}|_{{K}}\in\mathcal{P}_{1}(K)$, $\forall K\in\mathcal{A}_{h}$, then according to (36),
@@
$\mathrm{I}_{a}$ $\leq\sum_{K\in\mathcal{G}_{h}}||f-\nabla\cdot\bm{\sigma}_{h}||_{K\cap\Omega^{\star}}||v_{0}-I_{h}v_{0}||_{K\cap\Omega^{\star}}$
$\lesssim\Big{(}\sum_{K\in\mathcal{G}_{h}}h_{K}^{2}||f-\nabla\cdot\bm{\sigma}_{h}||^{2}_{K\cap\Omega^{\star}}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{K\in\mathcal{G}_{h}}h_{K}^{-2}||v_{0}-I_{h}v_{0}||^{2}_{K\cap\Omega^{\star}}\Big{)}^{\frac{1}{2}}$
$\leq\Big{(}\sum_{K\in\mathcal{G}_{h}}(\mathcal{E}_{\mathrm{div}}^{K})^{2}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{K\in\mathcal{T}_{h}^{0}}h_{K}^{-2}||v_{0}-I_{h}v_{0}||^{2}_{K}\Big{)}^{\frac{1}{2}}$
$\lesssim\Big{(}\sum_{K\in\mathcal{G}_{h}}(\mathcal{E}_{\mathrm{div}}^{K})^{2}\Big{)}^{\frac{1}{2}}||\nabla v_{0}||_{\Omega_{0}}\lesssim\Big{(}\sum_{K\in\mathcal{G}_{h}}(\mathcal{E}_{\mathrm{div}}^{K})^{2}\Big{)}^{\frac{1}{2}}||\nabla v||_{\Omega}.$ (51)
For $\mathrm{I}_{b}$ we have instead that, by the Cauchy–Schwarz inequality
$\mathrm{I}_{b}$ $\leq\sum_{K\in\mathcal{G}_{h}}||g+\bm{\sigma}_{h}\cdot\bm{n}||_{\gamma_{K}^{\star}}||v_{0}-I_{h}v_{0}||_{\gamma_{K}^{\star}}$
$\lesssim\Big{(}\sum_{K\in\mathcal{G}_{h}}h_{K}||g+\bm{\sigma}_{h}\cdot\bm{n}||^{2}_{\gamma_{K}^{\star}}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{K\in\mathcal{G}_{h}}h_{K}^{-1}||v_{0}-I_{h}v_{0}||^{2}_{\gamma_{K}^{\star}}\Big{)}^{\frac{1}{2}}$ (52)
As in [BCV_TRIMMING], let us recall the following local trace inequality, proven in [HH_2002] under the assumption of Lipschitz domains: there exists a fixed $h_{0}>0$ such that, if $K\in\mathcal{G}_{h}$ and $h_{K}<h_{0}$ then $\forall v\in H^{1}(K)$
$||v||^{2}_{\gamma_{K}^{\star}}$ $\lesssim h_{K}^{-1}||v||^{2}_{K}+h_{K}||\nabla v||^{2}_{K}.$ (53)
Hence, exploiting again (48)-(49), it follows that
$\sum_{K\in\mathcal{G}_{h}}h_{K}^{-1}||v_{0}-I_{h}v_{0}||_{\gamma_{K}^{\star}}^{2}$ $\lesssim\sum_{K\in\mathcal{G}_{h}}\big{(}h_{K}^{-2}||v_{0}-I_{h}v_{0}||_{K}^{2}+||\nabla(v_{0}-I_{h}v_{0})||_{K}^{2}\big{)}$
$\leq\sum_{K\in\mathcal{T}_{h}^{0}}\big{(}h_{K}^{-2}||v_{0}-I_{h}v_{0}||_{K}^{2}+||\nabla(v_{0}-I_{h}v_{0})||_{K}^{2}\big{)}$
$\lesssim\sum_{K\in\mathcal{T}_{h}^{0}}||\nabla v_{0}||_{K}^{2}=||\nabla v_{0}||_{\Omega_{0}}^{2}.$
Substituting in (52) and exploiting (46) we then have
$\mathrm{I}_{b}\lesssim\Big{(}\sum_{K\in\mathcal{G}_{h}}(\mathcal{E}_{\mathrm{g}}^{K})^{2}\Big{)}^{\frac{1}{2}}||\nabla v||_{\Omega}.$ (54)
For what concerns the term II in (47) we simply apply the Cauchy–Schwarz inequality, obtaining
$\mathrm{II}$ $\leq||\bm{\sigma}_{h}+\nabla u_{h}^{\star}||_{\Omega}||\nabla v_{0}||_{\Omega}\leq||\bm{\sigma}_{h}+\nabla u_{h}^{\star}||_{{\Omega^{\star}}}||\nabla v||_{\Omega}\leq\Big{(}\sum_{K\in\mathcal{A}_{h}}||\bm{\sigma}_{h}+\nabla u_{h}^{\star}||^{2}_{K\cap{\Omega^{\star}}}\Big{)}^{\frac{1}{2}}||\nabla v||_{\Omega},$
$=\Big{(}\sum_{K\in\mathcal{A}_{h}}(\mathcal{E}_{\sigma}^{K})^{2}\Big{)}^{\frac{1}{2}}||\nabla v||_{\Omega}$ (55)
@@
where we have used (41), and the fact that $\bm{n_{F}}=-\bm{n}$ Denoting by $\overline{v_{0}}^{\tilde{\gamma}_{F}}$ the average of $v_{0}$ over $\tilde{\gamma}_{F}$ and choosing $c=-\overline{v_{0}}^{\tilde{\gamma}_{F}}$, we can rewrite III as
@@
In order to bound $\mathrm{III}_{a}$ we use the Friedrichs inequality reported in [BCV2022_arxiv, Lemma A.2]: since $\tilde{\gamma}_{F}\subset\partial F$, $|\tilde{\gamma}_{F}|>0$ and $|\tilde{\gamma}_{F}|^{\frac{1}{d-1}}\simeq h_{F}$, then $\forall v\in H^{1}(F)$
$||v-\overline{v}^{\tilde{\gamma}_{F}}||_{F}\lesssim h_{F}||\nabla v||_{F}.$ (56)
We hence obtain, under Assumption 2 and using the Cauchy–Schwartz inequality, (56), (37) and (46), that
$\mathrm{III}_{a}$ $\leq\sum_{F\in\mathcal{\tilde{\mathcal{F}}}}||\nabla\cdot\bm{\sigma}_{h}-f||_{F}||v_{0}-\overline{v_{0}}^{\tilde{\gamma}_{F}}||_{F}\lesssim\sum_{F\in\mathcal{\tilde{\mathcal{F}}}}||\nabla\cdot\bm{\sigma}_{h}-f||_{F}\ h_{F}||\nabla v_{0}||_{F}$
$\leq\Big{(}\sum_{F\in\tilde{\mathcal{F}}}h_{F}^{2}||\nabla\cdot\bm{\sigma}_{h}-f||_{F}^{2}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{F\in\tilde{\mathcal{F}}}||\nabla v_{0}||_{F}^{2}\Big{)}^{\frac{1}{2}}$
$\lesssim\Big{(}\sum_{F\in\tilde{\mathcal{F}}}(h_{F}^{\min})^{2}\sum_{K\in\mathcal{G}_{h}}||\nabla\cdot\bm{\sigma}_{h}-f||_{F\cap K}^{2}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{F\in\tilde{\mathcal{F}}}||\nabla v_{0}||_{F}^{2}\Big{)}^{\frac{1}{2}}$
$\leq\Big{(}\sum_{F\in\tilde{\mathcal{F}}}\sum_{K\in\mathcal{G}_{h}}h_{K}^{2}||\nabla\cdot\bm{\sigma}_{h}-f||_{F\cap K}^{2}\Big{)}^{\frac{1}{2}}\Big{(}\sum_{F\in\tilde{\mathcal{F}}}||\nabla v_{0}||_{F}^{2}\Big{)}^{\frac{1}{2}}$
$\leq\Big{(}\sum_{K\in\mathcal{G}_{h}}h_{K}^{2}||\nabla\cdot\bm{\sigma}_{h}-f||_{K\cap{\Omega^{\star}}}^{2}\Big{)}^{\frac{1}{2}}||\nabla v_{0}||_{\Omega_{0}}\lesssim\Big{(}\sum_{K\in\mathcal{G}_{h}}(\mathcal{E}_{\mathrm{div}}^{K})^{2}\Big{)}^{\frac{1}{2}}||\nabla v||_{\Omega}.$ (57)
For what concerns $\mathrm{III}_{b}$ let us remark that $d_{F}=d_{F}^{h}+(\bm{\sigma}-\bm{\sigma}_{h})\cdot\bm{n}$ on $\tilde{\gamma}_{F}$, $\forall F\in\tilde{\mathcal{F}}$. We hence have that
@@
Once $\mathrm{III}_{b}$ is rewritten in this form, referring the reader to [BCV2022, Theorem 4.3] it is possible to prove that
$\mathrm{III}_{b}\lesssim\Big{(}\sum_{F\in\tilde{\mathcal{F}}}\mathcal{E}_{F}^{2}\Big{)}^{\frac{1}{2}}||\nabla v||_{\Omega},$ (58)
with $\mathcal{E}_{F}^{2}$ defined as in (42) and the hidden constant depending on Lipschitz, Poincaré and trace constants.
@@
$||\nabla e||_{\Omega}^{2}$ $\leq\Big{(}\sum_{K\in\mathcal{G}_{h}}\alpha_{1}(\mathcal{E}_{\mathrm{div}}^{K})^{2}\Big{)}^{\frac{1}{2}}||\nabla e||_{\Omega}+\Big{(}\sum_{K\in\mathcal{G}_{h}}\alpha_{2}(\mathcal{E}_{\mathrm{g}}^{K})^{2}\Big{)}^{\frac{1}{2}}||\nabla e||_{\Omega}+\Big{(}\sum_{K\in\mathcal{A}_{h}}(\mathcal{E}_{\sigma}^{K})^{2}\Big{)}^{\frac{1}{2}}||\nabla e||_{\Omega}$
$+\Big{(}\sum_{F\in\tilde{\mathcal{F}}}\alpha_{3}\mathcal{E}_{F}^{2}\Big{)}^{\frac{1}{2}}||\nabla e||_{\Omega}.$
@@
The defeaturing component of the estimator is very similar to the one proposed in [BCV2022, BCV2022_arxiv]. However, it is built from a reconstructed flux, which allows us not to evaluate the numerical flux $\nabla u_{h}^{\star}$ on the boundary of the features in $\tilde{\mathcal{F}}$. As observed also in [BCGVV], this is a great advantage in a finite element framework, in which the numerical flux is typically discontinuous across element edges.
@@ 5 The adaptive strategy
In the framework of adaptive finite elements for elliptic PDEs [NochettoPRIMER], at each iteration of the adaptive process we need to go though the following blocks:
@@
We now elaborate on each of these blocks, denoting by $s\in\mathbb{N}$ the current iteration index. We denote by $\Omega_{s}^{\star}$ the partially defeatured geometry at the beginning of the $s$-th iteration, choosing $\Omega_{0}^{\star}=\Omega_{0}$, and by $\mathcal{T}_{h}^{s}$ the mesh covering $\Omega_{0}$, obtained by the refinement of the starting mesh $\mathcal{T}_{h}^{0}$ through the $h$-adaptivity performed at the previous $s-1$ iterations. We remark again that the refined meshes are triangulations of $\Omega_{0}$, and therefore they are blind to the features.
@@ 5.1 Solve
At step $s$, a problem in the form of (16) is solved, with ${\Omega^{\star}}=\Omega_{s}^{\star}=\text{int}\big{(}\overline{\Omega}\cup\{\overline{F},~{}F\in\tilde{\mathcal{F}}_{s}\}\big{)}$, where ~$\tilde{\mathcal{F}}_{s}$ is the set of features still neglected (i.e. filled with material) at the beginning of iteration $s$, and which may be marked for geometric refinement at the end of the iteration. We define $\mathcal{F}_{s}^{\star}=\mathcal{F}\setminus\tilde{\mathcal{F}}_{s}$ as the set of features which have already been re-included into the geometry at the beginning of iteration $s$. The active portion $\mathcal{A}_{h}^{s}$ of the mesh $\mathcal{T}_{h}^{s}$ is defined as
@@ 5.2 Estimate
Let $\tilde{\gamma}_{0}=\tilde{\gamma}_{0,s}$, $\tilde{\gamma}=\tilde{\gamma}_{s}$, $\gamma_{0}^{\star}=\gamma_{0,s}^{\star}$ and ${\gamma^{\star}}=\gamma_{s}^{\star}$ be defined as in (11)-(12), with ~=~s~$\tilde{\mathcal{F}}=\tilde{\mathcal{F}}_{s}$ and $\mathcal{F}^{\star}=\mathcal{F}_{s}^{\star}$ and let
@@ 5.3 Mark
In order to mark the elements in $\mathcal{A}_{h}^{s}$ and the features in ~$\tilde{\mathcal{F}}_{s}$ we adopt a Dörfler strategy. Let us fix a marking parameter $\theta\in(0,1]$ and let $\mathcal{M}^{s}=\mathcal{A}_{h}^{s}\cup\tilde{\mathcal{F}}_{s}$. For each element or feature $M\in\mathcal{M}^{s}$ we define
@@ 5.4 Refine
and ~$\tilde{\mathcal{F}}_{s+1}=\tilde{\mathcal{F}}_{s}\setminus(\mathcal{D}^{s}\cap\tilde{\mathcal{F}}_{s})$, $\mathcal{F}_{s+1}^{\star}=\mathcal{F}_{s}^{\star}\cup(\mathcal{D}^{s}\cap\tilde{\mathcal{F}}_{s}).$ The active mesh for the following step is defined as $\mathcal{A}_{h}^{s+1}=\{K\in\mathcal{T}_{h}^{s+1}:~{}K\cap\Omega_{s+1}^{\star}\neq\emptyset\}$ (see Figure 5).
@@
It is interesting to remark how, similarly to [BCV2022_arxiv], the proposed adaptive strategy naturally takes care of Assumption 2. Let $F\in\tilde{\mathcal{F}}_{s}$, i.e. $F$ is a feature that is not included into the geometry at step $s$. If $F\cap K=\emptyset$ $\forall K\in\mathcal{G}_{h}^{s}$, which means that $F$ is not intersecting any element which is also intersected by a feature in $\mathcal{F}^{\star}_{s}$, then Assumption 2 is actually unnecessary for the given feature, since $||\nabla\cdot\bm{\sigma}_{h,s}-f||_{F}=0$ (see the proof of Proposition 1). However, it will still be naturally satisfied since, if the error is concentrated in the feature, this will lead to the inclusion of $F$ into the geometry. The same holds if there exists at least one element $K\in\mathcal{G}_{h}^{s}$ such that $K\cap F\neq\emptyset$. Indeed, the local contribution to the numerical component of the estimator in the elements in $\mathcal{G}_{h}^{s}$ is higher, due to the terms $\mathcal{E}_{\mathrm{div}}^{K}$ and $\mathcal{E}_{\mathrm{g}}^{K}$. This means that $K$ will be early marked for refinement, allowing to separate $F$ from the elements in $\mathcal{G}_{h}^{s}$ in a few iterations.

### v2

@@ 4 A posteriori error analysis
[…]
Let us remark that both (42) and (43) refer to the totally defeatured domain $\Omega_{0}$, so that the constants are completely independent from the position of the features with respect to the mesh.
Since ${I_{h}v_{0}}|_{{K}}\in\mathcal{P}_{r}(K)$, $\forall K\in\mathcal{A}_{h}$, then according to (30),
@@
[…]
with the last hidden constant depending on $C_{\mathrm{stein}}$ and $C_{\mathrm{pr,1}}$. For $\mathrm{I}_{b}$ we have instead that, by the Cauchy–Schwarz inequality
[…]
As in [9], let us recall the following local trace inequality, proven in [24] under the assumption of Lipschitz features: for every $K\in\mathcal{G}_{h}$, for $v\in H^{1}(K)$,
[…]
where $C_{\mathrm{tr}}$ is independent of $v$, of $K$, of how $\gamma^{\star}$ intersects $K$ and of $h<h_{0}$ for some arbitrary but fixed $h_{0}$. Hence, exploiting again (42)-(43), it follows that
[…]
with the last hidden constant depending on $C_{\mathrm{tr}}$, $C_{\mathrm{pr,1}}$ and $C_{\mathrm{pr,3}}$. Substituting in (46) and exploiting (40) we then have
[…]
where the hidden constant depends on $C_{\mathrm{tr}}$, $C_{\mathrm{pr,1}}$, $C_{\mathrm{pr,3}}$ and $C_{\mathrm{stein}}$. For what concerns the term II in (41) we simply apply the Cauchy–Schwarz inequality, obtaining
[…]
@@
where we have used (35), and the fact that $\bm{n_{F}}=-\bm{n}$ . Denoting by $\overline{v_{0}}^{\tilde{\gamma}_{F}}$ the average of $v_{0}$ over $\tilde{\gamma}_{F}$ and choosing $c=-\overline{v_{0}}^{\tilde{\gamma}_{F}}$, we can rewrite III as
@@
In order to bound $\mathrm{III}_{a}$ we use the Friedrichs inequality reported in [10, Lemma A.2]: since $\tilde{\gamma}_{F}\subset\partial F$, $|\tilde{\gamma}_{F}|>0$ and $|\tilde{\gamma}_{F}|^{\frac{1}{d-1}}\simeq h_{F}$, then $\forall v\in H^{1}(F)$
[…]
where $C_{\mathrm{f}}$ is independent of the size of $F$. We hence obtain, under Assumption 2 and using the Cauchy–Schwartz inequality, (50), (31) and (40), that
[…]
with the last hidden constant depending on $C_{\mathrm{f}}$, $C_{\mathrm{stein}}$ and the constant in Assumption 2. For what concerns $\mathrm{III}_{b}$ let us remark that $d_{F}=d_{F}^{h}+(\bm{\sigma}-\bm{\sigma}_{h})\cdot\bm{n}$ on $\tilde{\gamma}_{F}$, $\forall F\in\tilde{\mathcal{F}}$. We hence have that
@@
Once $\mathrm{III}_{b}$ is rewritten in this form, referring the reader to [8, Theorem 4.3] it is possible to prove that
[…]
with $\mathcal{E}_{F}^{2}$ defined as in (36) and the hidden constant depending on the Poincaré constant on $\tilde{\gamma}_{F}$ (see [8, Appendix A.1]) and a trace constant on $\Omega$ .
@@
[…]
@@
Remark 5.
An eligible candidate for operator $I_{h}$ in (42)-(43) is a Scott-Zhang quasi-interpolation operator, mapping $v_{0}\in H_{0,\Gamma_{\mathrm{D}}}^{1}(\Omega_{0})$ into $I_{h}v_{0}\in V_{h}^{0}(\Omega_{0})\subset Q_{h}(\Omega_{0})$. However, as (30) allows for discontinuous test functions, an $L^{2}$-projection can also be considered, possibly leading to lower hidden constants (see e.g. [6]).
@@
The defeaturing component of the estimator is very similar to the one proposed in [8, 10]. However, it is built from a reconstructed flux, which allows us not to evaluate the numerical flux $\nabla u_{h}^{\star}$ on the boundary of the features in $\tilde{\mathcal{F}}$. As observed also in [7], this is a great advantage in a finite element framework, in which the numerical flux is typically discontinuous across element faces. Indeed, the feature boundary can be arbitrarily close to mesh faces, so that evaluating the numerical flux along this line may result in spurious jumps. This issue is avoided by using an equilibrated flux reconstruction, which is uniquely defined on element faces.
@@
Remark 8.
Let us observe that the constants in Proposition 1 are independent of how the feature boundaries intersect the mesh. Indeed, badly cut elements do not directly cause these constants to blow up; rather, their impact arises from the potential loss of stability of the flux reconstruction.
@@ 5 The adaptive strategy
In the framework of adaptive finite elements for elliptic PDEs [30], at each iteration of the adaptive process we need to go though the following blocks:
@@
We now elaborate on each of these blocks, denoting by $s\in\mathbb{N}$ the current iteration index. We denote by $\Omega_{s}^{\star}$ the partially defeatured geometry at the beginning of the $s$-th iteration, choosing $\Omega_{0}^{\star}=\Omega_{0}$, and by $\mathcal{T}_{h}^{s}$ the mesh covering $\Omega_{0}$, obtained by the refinement of the starting mesh $\mathcal{T}_{h}^{0}$ through the $h$-adaptivity performed at the previous s1s-1 iterations. We remark again that the refined meshes are triangulations of $\Omega_{0}$, and therefore they are blind to the features.
@@ 5.1 Solve
At step $s$, a problem in the form of (16) is solved, with ${\Omega^{\star}}=\Omega_{s}^{\star}=\text{int}\big(\overline{\Omega}\cup\{\overline{F},~F\in\tilde{\mathcal{F}}_{s}\}\big)$, where $\tilde{\mathcal{F}}_{s}$ is the set of features still neglected (i.e. filled with material) at the beginning of iteration $s$, and which may be marked for geometric refinement at the end of the iteration. We define $\mathcal{F}_{s}^{\star}=\mathcal{F}\setminus\tilde{\mathcal{F}}_{s}$ as the set of features which have already been re-included into the geometry at the beginning of iteration $s$. The active portion $\mathcal{A}_{h}^{s}$ of the mesh $\mathcal{T}_{h}^{s}$ is defined as
@@ 5.2 Estimate
Let $\tilde{\gamma}_{0}=\tilde{\gamma}_{0,s}$, $\tilde{\gamma}=\tilde{\gamma}_{s}$, $\gamma_{0}^{\star}=\gamma_{0,s}^{\star}$ and ${\gamma^{\star}}=\gamma_{s}^{\star}$ be defined as in (11)-(12), with $\tilde{\mathcal{F}}=\tilde{\mathcal{F}}_{s}$ and $\mathcal{F}^{\star}=\mathcal{F}_{s}^{\star}$ and let
@@ 5.3 Mark
In order to mark the elements in $\mathcal{A}_{h}^{s}$ and the features in $\tilde{\mathcal{F}}_{s}$ we adopt a Dörfler strategy [16]. Let us fix a marking parameter $\theta\in(0,1]$ and let $\mathcal{M}^{s}=\mathcal{A}_{h}^{s}\cup\tilde{\mathcal{F}}_{s}$. For each element or feature $M\in\mathcal{M}^{s}$ we define
@@ 5.4 Refine
and $\tilde{\mathcal{F}}_{s+1}=\tilde{\mathcal{F}}_{s}\setminus(\mathcal{D}^{s}\cap\tilde{\mathcal{F}}_{s})$, $\mathcal{F}_{s+1}^{\star}=\mathcal{F}_{s}^{\star}\cup(\mathcal{D}^{s}\cap\tilde{\mathcal{F}}_{s}).$ The active mesh for the following step is defined as $\mathcal{A}_{h}^{s+1}=\{K\in\mathcal{T}_{h}^{s+1}:~K\cap\Omega_{s+1}^{\star}\neq\emptyset\}$ (see Figure 5).
@@
It is interesting to remark how, similarly to [10], the proposed adaptive strategy naturally takes care of Assumption 2. Let $F\in\tilde{\mathcal{F}}_{s}$, i.e. $F$ is a feature that is not included into the geometry at step $s$. If $F\cap K=\emptyset$ $\forall K\in\mathcal{G}_{h}^{s}$, which means that $F$ is not intersecting any element which is also intersected by a feature in $\mathcal{F}^{\star}_{s}$, then Assumption 2 is actually unnecessary for the given feature, since $\nabla\cdot\bm{\sigma}_{h,s}-f||_{F}=0$ (see the proof of Proposition 1). However, it will still be naturally satisfied since, if the error is concentrated in the feature, this will lead to the inclusion of $F$ into the geometry. The same holds if there exists at least one element $K\in\mathcal{G}_{h}^{s}$ such that $K\cap F\neq\emptyset$. Indeed, the local contribution to the numerical component of the estimator in the elements in $\mathcal{G}_{h}^{s}$ is higher, due to the terms $\mathcal{E}_{\mathrm{div}}^{K}$ and $\mathcal{E}_{\mathrm{g}}^{K}$. This means that $K$ will be early marked for refinement, allowing to separate $F$ from the elements in $\mathcal{G}_{h}^{s}$ in a few iterations.